A pizza parlour will deliver a pizza take-away order to the customer if s/he lives within a 5 kilometre radius from the pizza parlour. If the customer lives within this radius, it is found that the time taken to receive the pizza is normally distributed with a mean time of 45 minutes and a standard deviation of 8 minutes. Paul has relocated and now lives within the 5 kilometre radius. He orders a pizza from the pizza parlour. The probability that it takes between 43 and 49 minutes for him to receive his pizza is


1
Expert's answer
2021-10-12T14:09:03-0400

Let "X=" the time taken to receive the pizza: "X\sim N(\mu, \sigma^2)."

Given "\mu=45\ min, \sigma=8\ min."

"P(43<X<49)=P(X<49)-P(X\leq 43)"

"=P(Z< \dfrac{49-\mu}{\sigma})-P(Z\leq \dfrac{43-\mu}{\sigma})"

"=P(Z< \dfrac{49-45}{8})-P(Z\leq \dfrac{43-45}{8})"

"=P(Z< 0.5)-P(Z\leq-0.25)"

"\approx0.69146246-0.40129367"

"\approx0.290169"


The probability that it takes between 43 and 49 minutes for him to receive his pizza is "0.290169."